OpenStudy (anonymous):

The volume of two similar solids is 1331 m3 and 729 m3. The surface area of the larger solid is 605 m2. What is the surface area, in square meters, of the smaller solid? Answer 81 121 305 405

OpenStudy (anonymous):

Is a square, cuboid or what ??

OpenStudy (anonymous):

yes, a square is a cuboid

OpenStudy (phi):

the solids are similar so you know that their sides are in proportion \[\frac{\sqrt[3]{1331}} { \sqrt[3]{729}}\] their surface areas will be in the ratio of the squares of the side.

OpenStudy (anonymous):

okk thank-you!